OpenStudy (anonymous):

Name the property of real numbers illistrated by the equation 7(x-y)=7x-7y

OpenStudy (anonymous):

multiplicative

hero (hero):

Distributive property a(b + c) = ab + ac 7(x - y) = 7x - 7y Same structure